The properties of addition are:
[1] Commutative Property
[2] Associative Property
[3] Additive Identity Property
[4] Additive Inverse Property
For each equation below, indicate the property that justifies the equation by filling in the box with the appropriate number.
b=b+0
7+(6+y)=(7+6)+y
0=x+(-x)

Answer by stanbon(75887)

The properties of addition are:
[1] Commutative Property
[2] Associative Property
[3] Additive Identity Property
[4] Additive Inverse Property
For each equation below, indicate the property that justifies the equation by filling in the box with the appropriate number.
-----------------
b=b+0
Ans: #3
---------------
7+(6+y)=(7+6)+y
Ans: #2
---------------
0=x+(-x)
Ans: 4
===========================
Cheers,
Stan H.